Solve the system below. Parameterize the solution set 10w + 2z 5y -3 16w 4Z 5y + 11 2u 33w + 7z 15у + 3x у+
Answer #1

Given ststem of equation is 5tat-10w u = 6 - - -3 -5t 4t-16w 3x -157 +72 -33w +2u 1 form Matrin -5 3 5 -16 -3 2 -15 -33 A AxAppl R32 R3 - 5 -1 2. -10 -6 2. - -6-2-14 2 Appl R R-R3 5 2. 6 2. 0 Trcanforem into equoti 0 Forcm 6 -5オ+22-IDW tu 2 -6w -u ニfreom ea ti) 22 -6W - しニー9 =そてと -9 +6w +Uu -4 t6w 1 9 +6w+ 5 = -2+3w from eg i -57+22-10w+uo -X = 6+5オー22ナ10w-u - 2 +3w-+1o-